OpenStudy (doodleduck8):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. y-1=6(x+5) y+5=6(x-1) y-5=6(x+1) y+1=6(x-5)

OpenStudy (doodleduck8):

@shirleyarnett

OpenStudy (will.h):

Use Y - y1 = m(x - x1) M is the slope And substitute Y + 1 = 6(x - 5)